A concentrated HCl solution contains 37% w/w HCl. The specific gravity of the solution is 1.19 g/mL. Calculate its molarity
1
Expert's answer
2020-11-26T10:55:32-0500

CM=n/V

100 g of a given solution has a volume of:

V (HCl) = 100/1.19 = 84.03 ml

n = m/M

M (HCl) = 36.5 mol/l

n (HCl) = 37/36.5 = 1.01 mol

CM (HCl) = 1.01/84.03 = 0.01 mol/l